• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Chai

    Log entry time 08:03:17 on June18,2000

    Entry number 44118

    keyword=Owl shift summary



    00:00 Owl shift starts.
    Still no beam. MCC is still trying to fix the vacuum problem in north linac.
    It may take 2 more hours.
    01:44 MCC said they fixed the problem in the linac and is going to tune the beam.
    It will take about 30 min.
    02:45 Beam back to Hall A (71.6uA).
    Continue taking data at kin(135,180).

    time run events comment
    02:45 2196 0.6M PS1 and PS3 are 1.
    02:49 2197 1.3M
    03:17 2198 1.3M
    03:44 2199 1.3M
    04:11 2200 1.3M
    04:37 2201 1.3M
    05:02 2202 1.3M
    05:29 2203 1.3M
    05:57 2204 1.3M
    06:27 2205 1.3M
    06:53 2206 1.3M
    07:19 2207 1.3M
    07:46 2208